# Superfluid Dark Matter

### v1

@@ 4.2 Caveats of structure formation
Alternatively, instead of a coherent state, DM can be in the form of an ultra-cold gas of massive particles. While these two possibilities are indistinguishable at the background level, the evolution of perturbations might be different. In this scenario, since we are dealing with a collection of weakly interacting particles, the DM particles are expected to acquire a velocity dispersion determined by the virial theorem. In reality, without a concrete UV completion of our scalar field model (unlike axions) and an explicit production mechanism, the initial state should be considered a mixture of both coherent and gaseous components.
@@
It must be stressed that our primary interest here is not the classical growth of perturbations but rather the deviation from coherence as it occurs. Therefore, the c-number functions $\langle\hat{\Psi}\rangle$ and $\Phi_{\text{\tiny N}}$ in Eq. (4.13) are assumed to be solutions of the classical equations of motion through the non-linear process of structure formation/virialization.303030 Depending on the scales of interest, Eq. (4.13) can be simplified accordingly. For instance, noting that $|\langle\hat{\Psi}\rangle|^{2}\sim\rho/m$ and $\Phi_{\text{\tiny N}}\sim G\rho k^{-2}$, it is straightforward to see that the last term in Eq. (4.13) can be neglected on scales significantly shorter than the Jeans scale. This is indeed a valid approximation, since we are interested in the plausibility of the production of particles with de Broglie wavelength set by the characteristic CDM velocity, and thus possibly smaller than the Jeans scale. Solving for the mode functions of $\hat{\psi}$ in the background of $\Phi_{\text{\tiny N}}$ and $\langle\hat{\Psi}\rangle$ would allow us to calculate the tree-level correlation functions, which in turn can be used to calculate the depleting backreaction on the one-point function itself.
@@
In the context of fuzzy DM, quantum corrections to the coherent classical evolution have been numerically studied in Refs. [210, 211] in the non-relativistic approximation. There, the authors studied the spread of the wavefunction around its peak value, and found that there is an exponential growth in configurations undergoing a classical non-linear evolution. In contrast, the growth is only quadratic in time for stable, virialized systems.  [^31] This agrees with the general statement of Ref. [204] which shows that, for systems exhibiting classical (or Lyapunov) instabilities, the timescale over which classical evolution remains valid scales logarithmically with the number of constituent particles of the system—a large number for coherent configurations. In contrast, for stable systems, this timescale follows a power-law scaling. As a result, the breakdown of classicality occurs much more rapidly in the presence of classical instabilities [204, 205]. These numerical studies conclude that the loss of coherence is always a subleading effect in fuzzy DM. In the context of SDM, on the other hand, the different mass range and the role of self-interactions may play a more important role in the way the loss of coherence is realized. Therefore, further investigations are essential to carry out this computation in details, ensuring the validity of this framework and refining our understanding of the cosmological evolution of the SDM candidate.
@@ 5.1 Merging galaxies: Bullet Cluster
The upshot of Eqs. (5.1) and (5.2) is that the Bullet Cluster system effectively places a constraint on the interaction rate $\Gamma$. The resulting constraint on the model parameters depends on the specific DM theory under consideration, but also on the phase space distribution of DM within halos. [^33] Let us mention that the approximation underlying Eq. (5.2) is valid, in the context of SDM, provided that the typical involved velocities are larger than the sound speed of the superfluid. In particular, in this limit, the collision of two superfluids can be approximated as the collision of on-shell constituent particles. See also Appendix A.
@@ 5.2 Rotation curves
Nonetheless, there are still cases where we can make some qualitative predictions about the halo density distribution. In particular, while the dynamics of larger galactic halos, such as that of the Milky Way, are expected to be dominated by the out-of-equilibrium and debris components, this is not the case for dwarf galaxies. By adjusting the parameters of the quartic superfluid theory, it is possible to have a central core extending approximately up to $10-15$ kpc from the galactic center, thereby engulfing the majority of the DM fraction of dwarf galaxies, according to the relation
@@
Before concluding this Section, it is worth highlighting another key avenue for testing ultralight DM models: the scaling relation between core density and core radius in galaxies. Not only should these models predict the existence of galactic cores, but they must also consistently account for the cores observed across many—if not all—galaxies. While model parameters can be adjusted to fit the core of a single galaxy, a robust model must accurately reproduce the observed scaling relation across a broader population of galaxies. This issue has been investigated both in the context of fuzzy and superfluid DM models [221, 222, 223, 224, 197]. For the latter, recent hydrodynamical simulations indicate that the scaling relations of simulated SDM halos fail to match those of Milky Way dwarf spheroidals and nearby galaxies—even for kpc-scale core radii—thereby falling short of resolving the cusp-core and too-big-to-fail problems simultaneously [197]. A thorough understanding of the SDM scaling relations remains an open and active area of research. In particular, incorporating baryonic physics is necessary for a more definitive comparison with data. Improvements along these directions are left to future work.
@@ 5.3 Vortices
×$\vec{\nabla}\times\vec{v}=\frac{1}{m}2\pi\delta^{(3)}\big{(}\vec{r}-\vec{r}_{0}\big{)}\hat{z}\,,$ (5.13)
@@
where $r$ is the distance from the vortex, $\varphi$ is the polar angle, and $n$ is the integer winding number. [^36] The winding number should not be mistaken for the number density of Eq. (2.7). The function $f(r)$ determines the profile of the configuration that interpolates from the vortex line to the asymptotically unperturbed superfluid (i.e.$f(r\rightarrow\infty)=1$). Substituting the above ansatz into (2.30) gives the equation for $f$:
@@
$L_{z}=n\mu v^{2}\int{\rm d}^{3}xf^{2}\simeq\pi n\frac{\rho}{m}R^{3}\cdot\begin{cases}1&~{}~{}~{}~{}\textit{asymptotics}\,;\\ (n+1)^{-1}(2mc_{s}R)^{-2}&~{}~{}~{}~{}\textit{vortex throat}\,,\end{cases}$ (5.18)
@@
This naturally brings us to the point of vortex formation in our SDM halos, which may have some nonzero angular momentum, [^38] The presence of angular momentum in scalar DM models can also arise in an inhomogeneous density field, or if the patch has a nonspherical boundary, such as in Riemann S-ellipsoids [228]. see Refs. [231, 232, 233, 173, 229] for related works. Indeed, halos tend to form with rotation angular frequency parameterized as $\Omega\simeq\chi_{\rm s}\sqrt{G\rho}$, where the spin parameter $\chi_{\rm s}$ takes values in the range $0.01\lesssim\chi_{\rm s}\lesssim 0.1$, as inferred from N-body simulations [234].

### v2

@@ 4.2 Caveats of structure formation
Alternatively, instead of a coherent state, DM can be in the form of an ultra-cold gas of massive particles. While these two possibilities are indistinguishable at the background level, the evolution of perturbations might be different. In this scenario, since we are dealing with a collection of weakly interacting particles, the DM particles are expected to acquire a velocity dispersion determined by the virial theorem. In reality, without a concrete UV completion of the scalar field model (unlike axions) and an explicit production mechanism, the initial state should be considered a mixture of both coherent and gaseous components.
@@
It must be stressed that our primary interest here is not the classical growth of perturbations but rather the deviation from coherence as it occurs. Therefore, the c-number functions $\langle\hat{\Psi}\rangle$ and $\Phi_{\text{\tiny N}}$ in Eq. (4.13) are assumed to be solutions of the classical equations of motion through the non-linear process of structure formation/virialization. [^35] Depending on the scales of interest, Eq. (4.13) can be simplified accordingly. For instance, noting that $|\langle\hat{\Psi}\rangle|^{2}\sim\rho/m$ and $\Phi_{\text{\tiny N}}\sim G\rho k^{-2}$, it is straightforward to see that the last term in Eq. (4.13) can be neglected on scales significantly shorter than the Jeans scale. This is indeed a valid approximation, since we are interested in the plausibility of the production of particles with de Broglie wavelength set by the characteristic CDM velocity, and thus possibly smaller than the Jeans scale. Solving for the mode functions of $\hat{\psi}$ in the background of $\Phi_{\text{\tiny N}}$ and $\langle\hat{\Psi}\rangle$ would allow us to calculate the tree-level correlation functions, which in turn can be used to calculate the depleting backreaction on the one-point function itself.
@@
In the context of fuzzy DM, quantum corrections to the coherent classical evolution have been numerically studied in Refs. [175, 176] in the non-relativistic approximation. There, the authors studied the spread of the wavefunction around its peak value, and found that there is an exponential growth in configurations undergoing a classical non-linear evolution. In contrast, the growth is only quadratic in time for stable, virialized systems.  [^36] This agrees with the general statement of Ref. [169] which shows that, for systems exhibiting classical (or Lyapunov) instabilities, the timescale over which classical evolution remains valid scales logarithmically with the number of constituent particles of the system—a large number for coherent configurations. In contrast, for stable systems, this timescale follows a power-law scaling. As a result, the breakdown of classicality occurs much more rapidly in the presence of classical instabilities [169, 171]. These numerical studies conclude that the loss of coherence is always a subleading effect in fuzzy DM. In the context of SDM, on the other hand, the different mass range and the role of self-interactions may play a more important role in the way the loss of coherence is realized. Therefore, further investigations are essential to carry out this computation in details, ensuring the validity of this framework and refining the understanding of the cosmological evolution of the SDM candidate.
@@ 5.1 Merging galaxies: Bullet Cluster
The upshot of Eqs. (5.1) and (5.2) is that the Bullet Cluster system effectively places a constraint on the interaction rate $\Gamma$. The resulting constraint on the model parameters depends on the specific DM theory under consideration, but also on the phase space distribution of DM within halos.
@@
Let us stress that the approximation underlying Eq. (5.2) is valid in the context of SDM provided that the infall velocity exceeds the sound speed of the superfluid. In this regime, the collision between the two superfluid halos can be approximated as a head-on collision of on-shell constituent particles, which is also the standard approximation typically adopted for CDM. If this condition is not met, the correct degrees of freedom describing the collision are phonons with an almost linear dispersion relation. In this case, the dynamics of the process can no longer be described as the scattering of individual constituents; rather, it must be analyzed in terms of interactions between the background fields, the resulting collective (and soft) phonon excitations, and potentially significant surface effects. However, in SDM, the sound speed is constrained to remain below $c_{s}\sim 10^{-3}$ in order to satisfy the bound from matter–radiation equality (4.6), so it is at least one order of magnitude smaller than the infall velocity. This justifies the approximation employed in this Section of considering the 2-body scattering of constituents as the leading interaction channel between the solitons. A detailed discussion of why the scattering between superfluids with relative velocities much greater than the sound speed reduces to the standard constituent scattering is presented in Appendix A.
@@ 5.2 Rotation curves
Nonetheless, there are still cases where we can make some qualitative predictions about the halo density distribution. In particular, while the dynamics of larger galactic halos, such as that of the Milky Way, are expected to be dominated by the out-of-equilibrium and debris components, this is not the case for dwarf galaxies. By adjusting the parameters of the quartic superfluid theory, it is possible to have a central core extending approximately up to 101510-15 kpc from the galactic center, thereby engulfing the majority of the DM fraction of dwarf galaxies, according to the relation
@@
5.3 Structure and dynamics of superfluid dark matter halos
In this Section, we present an overview of the qualitative expectations for the structure and dynamics of DM halos within the SDM framework, based on the analytic considerations developed thus far. Firm, quantitative predictions will ultimately require dedicated numerical simulations—which we leave for future work—especially to capture the intricate dynamics of tidal disruption within halos and the subsequent evolution of the superfluid debris. Nevertheless, the existing analytic treatment already points toward several generic features of the model. In particular, cores are naturally formed in SDM due to the formation of central solitons, providing a potential resolution to the cusp-core problem observed in galaxies. We outline the expected scaling behaviors, discuss the tendency toward cored versus cuspy density profiles, and comment on how SDM may account for the observed diversity of galactic rotation curves. These considerations sketch the main phenomenological implications of SDM on galactic scales and help guide the direction of future, simulation-based studies.
Following the detailed discussion in Sec. 3, the main expectation for the central regions of DM halos in the SDM framework is the formation of a central soliton, supported by the balance between repulsive self-interactions and gravity. The size and mass of this core depend sensitively on the microscopic theory—in particular on the equation of state and the density-dependence of the sound speed $c_{s}(\rho)$, which can be nontrivial in models beyond the simple quartic case (where $c_{s}$ depends linearly on the density). Surrounding the soliton, the superfluid phase is generically prone to gravitational fragmentation, governed by the Jeans scale of the condensate. For a superfluid with density $\rho$, the Jeans scale reads from Eq. (3.6), $\lambda_{\rm J}\simeq\sqrt{\frac{\pi c_{s}^{2}}{G\rho}}$, which also sets a characteristic minimum mass for the halo, $M_{\text{\tiny DM halo}}\gtrsim\frac{4\pi}{3}\,\rho\,\lambda_{\rm J}^{3}$. This implies that halos capable of hosting a central soliton cannot be lighter than this quantity, which depends sensitively on the microphysical parameters of the theory. Most of the resulting droplets from fragmentation are expected to experience strong tidal disruption, dissolving into extended streams of superfluid debris rather than surviving as long-lived bound clumps. Capturing this nonlinear, multi-scale dynamics— especially the shredding, mixing, and re-condensation of tidally disrupted fragments—requires numerical simulations.
We also emphasize that outliers may exist: in particular, regions near the outskirts of halos may fail to form a coherent soliton, and satellite halos may instead originate from the condensation of tidal superfluid streams far from the central core, with masses smaller than $\frac{4\pi}{3}\,\rho\,\lambda_{\rm J}^{3}$. Although these analytic considerations outline a qualitative picture—a central soliton, a fragmented superfluid, and debris-dominated outskirts—the detailed structure of halos in SDM ultimately awaits numerical simulations capable of resolving tidal disruption, phase transitions, and the evolution of superfluid streams, both within the central and far regions of DM halos. This preliminary understanding, however, allows us to begin addressing a few of the common puzzles and features that characterize CDM, as well as other DM candidates such as fuzzy DM, in terms of halo structure, core formation, and diversity of galactic rotation curves.
Let us start by discussing the diversity problem in galactic rotation curves. In standard CDM, galaxies with similar total halo masses are predicted to host relatively self-similar, cuspy density profiles [363, 362, 404, 394, 385], which implies a narrow range of circular velocities at fixed outer halo velocity. Observations, however, show a much wider spread in inner rotation curves, indicating that CDM alone may not fully capture the processes shaping the inner halo structure. Baryonic feedback can partially modify central densities, but it is likely insufficient to explain the full observed diversity, particularly in low-mass or DM-dominated systems [404]. In models of ultralight DM, such as fuzzy DM, one expects the presence of central solitonic cores embedded in a larger halo. In this theory, simulations show that the core-halo relation exhibits substantial scatter [122, 405], primarily due to the ambiguous transition between the solitonic core and the outer NFW-like halo, as well as due to the halo’s merger history, tidal interactions, and relaxation state. Despite this, halos always host a central core, and including baryonic effects may further modify the core-halo relation. This highlights that the inner halo structure in fuzzy DM is highly sensitive to both microphysics and dynamical history, offering a useful point of comparison for SDM, where similar effects are likewise expected to play a role. As we discussed in Sec. 5.2 on rotation curves, the presence of baryons in halos could similarly modify the core properties in SDM, as illustrated in Fig. 6. Just as in fuzzy DM, where the core-halo structure is sensitive to merger history, tidal interactions, and relaxation state, the inner structure of SDM halos can be affected by both the superfluid microphysics and baryonic processes. This analogy suggests that the diversity of inner halo properties, including core size and density, may naturally arise in SDM, providing a framework to interpret the range of observed galactic rotation curves.
@@ 5.3 Structure and dynamics of superfluid dark matter halos
A further aspect worth considering is the scaling relation between core density and core radius in galaxies. Not only should these models predict the existence of galactic cores, but they must also consistently account for the cores observed across many—if not all—galaxies. While model parameters can be adjusted to fit the core of a single galaxy, a robust model must accurately reproduce the observed scaling relation across a broader population of galaxies. In the case of the quartic superfluid theory, the prediction for the core would appear as roughly a vertical line in a plot of core density versus core radius. However, in general it is not straightforward to clearly identify the core, due to the presence of superfluid debris and fragmented structures surrounding the central region. The general problem of reproducing the observed core density–radius relation has been investigated both in the context of fuzzy and superfluid DM models [150, 104, 380, 51, 228, 78]. For the latter, recent hydrodynamical simulations indicate that the scaling relations of simulated SDM halos fail to match those of Milky Way dwarf spheroidals and nearby galaxies—even for kpc-scale core radii—thereby falling short of resolving the cusp-core and too-big-to-fail problems simultaneously [228]. A thorough understanding of the SDM scaling relations remains an open and active area of research. In particular, incorporating baryonic physics is necessary for a more definitive comparison with data. Improvements along these directions are left to future work.
As a final point, we comment on ultra‑faint dwarf galaxies, the smallest and most DM‑dominated galaxies known. Their stellar half‑light radii are only of order tens of parsecs (often $\lesssim 50$ pc), and their velocity dispersions are extremely small, implying mass‑to-light ratios among the highest observed. Because of their low stellar content and high DM densities, ultra‑faint dwarf galaxies provide a nearly “clean” laboratory for testing wave‑like (ultra‑light) candidates with minimal baryonic complications. They have been used to place informative constraints on the particle mass of fuzzy DM, mainly due to dynamical heating from wave interference in the DM, which is mostly set by the de-Broglie wavelength of the particles [137, 331, 460, 174]. Additional effects, such as tidal forces from the Milky Way, stochastic interference, soliton fluctuations, and tidal stripping, can modify the expected heating of stars and must be carefully considered when interpreting constraints from these systems [137, 331, 460, 174].
Let us now discuss the formation of these galaxies in the context of SDM. As shown above, the central soliton typically has a size of order kpc for certain choices of the microscopic parameters, such as the particle mass and self-interaction strength. One possible mechanism to form bounded, core‑like satellite structures is through the superfluid debris surrounding the central soliton. For such debris, the coherence scale is not simply given by the Jeans scale of Eq. (3.6), since the assumption of a self-gravitating, spherically symmetric structure is broken (e.g., due to tidal stripping from the host halo). In general, the relevant coherence scale of the debris is expected to be smaller, likely set by the tidal radius of the forming satellite, giving rise to potentially small structures, $\lambda_{\rm J}^{\text{\tiny satellites}}\ll\lambda_{\rm J}$. If the mini halo forms instead in the field, rather than as a satellite of a larger system, we expect a genuine central core to develop. However, its size may be relatively small because the central density in such isolated mini halos is typically low, resulting in correspondingly small core masses. Once such satellite or field structures form, it is unclear whether tidal heating occurs in the same way as in fuzzy DM. In SDM, the effective de-Broglie wavelength is much shorter because the particle mass is significantly larger than in fuzzy DM, and typical virial velocities are similar, thus potentially alleviating the effects on stars. Moreover, self‑interactions can further modify the dynamics. Consequently, more detailed studies and dedicated simulations are needed to assess whether bounds analogous to those derived from ultra‑faint dwarf galaxies in fuzzy DM can be established for SDM, and if so, how they differ in magnitude and scaling.
Further work is needed to fully address these issues, particularly through numerical simulations capable of resolving tidal disruption, debris evolution, and core formation in diverse halo environments.
5.4 Portals to Standard Model
In the foregoing Sections, we have presented a model in which DM is constituted by a self-interacting scalar field, whose dynamics and phenomenology are determined entirely within the dark sector. Importantly, in its simplest incarnation, the model is minimal in the sense that:
  • There are no a priori couplings linking the dark sector to the SM. In particular, we do not assume any portal interactions (e.g., Higgs, vector, or axion-like portals);
  • The dark sector evolution proceeds purely through its self-interactions and gravitational coupling to the SM.
Because of this minimal construction, all conventional bounds that rely on energy transfer to SM particles are automatically evaded. In this sense, the dark fluid can be viewed as a self-contained and gravitationally coupled component of the universe.
However, the absence of SM couplings is a choice of minimality, not a fundamental requirement. If one were to introduce portal interactions, the model would naturally fall into the broader class of scalar or axion-like DM scenarios, though it differs in key aspects—most notably by exhibiting stronger self-interactions. In that case, the model can be subjected to associated astrophysical and laboratory constraints.
To draw an analogy with alternative DM model in the same ultralight mass range, one can consider the QCD axion [454, 457], as briefly discussed in Sec. 1.2. The QCD axion is the pseudo Nambu–Goldstone boson $a$ of a spontaneously broken Peccei–Quinn (PQ) symmetry [375, 374], with potential arising from non-perturbative effects of the form [441]
$V(a)=m_{a}^{2}f_{a}^{2}\left[1-\cos\left(\frac{a}{f_{a}}\right)\right]\,$, (5.12)
in terms of the axion mass $m_{a}$ and PQ symmetry-breaking scale $f_{a}$, related by the condition [216]
$m_{a}\simeq 5.7~\mu\mathrm{eV}\left(\frac{10^{12}\,\mathrm{GeV}}{f_{a}}\right)\,.$ (5.13)
By expanding the potential for small field amplitudes, one gets
$V(a)=\frac{1}{2}m_{a}^{2}a^{2}-\frac{1}{4!}\frac{m_{a}^{2}}{f_{a}^{2}}a^{4}+\mathcal{O}(a^{6})\,$, (5.14)
which corresponds to an attractive quartic self-interaction
$\lambda_{4}=-\frac{m_{a}^{2}}{6f_{a}^{2}}\,.$ (5.15)
Equation (5.15) highlights a striking contrast between superfluid and axion DM: while the stability of the former relies on strong repulsive self-interactions that counteract gravity, the latter features much weaker ($m_{a}\ll f_{a}$) attractive self-interactions that instead enhance gravitational collapse.
The QCD axion can couple to the SM fields (including photons, electrons and neutrons) through higher-dimensional operators of the form (see [157] for a recent review)
$\mathcal{L}_{\text{\tiny int}}\supset\frac{g_{a\gamma\gamma}}{4}\,aF_{\mu\nu}\tilde{F}^{\mu\nu}+{\rm i}g_{aee}\,\partial_{\mu}a\,\bar{e}\gamma^{\mu}\gamma_{5}e+{\rm i}g_{aNN}\,\partial_{\mu}a\,\bar{N}\gamma^{\mu}\gamma_{5}N+\cdots\,$, (5.16)
with typical magnitudes (in terms of $\mathcal{O}(1)$ coefficients)
$g_{a\gamma\gamma}\sim\frac{\alpha}{2\pi f_{a}}\,,\qquad g_{aee}\sim C_{e}\frac{m_{e}}{f_{a}}\,,\qquad g_{aNN}\sim C_{N}\frac{m_{N}}{f_{a}}\,.$ (5.17)
These couplings are extremely small, typically in the range $g_{a\gamma\gamma}\sim(10^{-13}\div 10^{-10})\,\mathrm{GeV}^{-1}$, but potentially not vanishing, allowing the axion to interact weakly with SM particles while remaining a viable DM candidate [5]. Let us emphasize that an analogous discussion applies to ALPs, which share many features with QCD axions, including similar EFTs and a potential role as DM. However, unlike QCD axions, ALPs do not couple to gluons in the same way and, since their masses are not generated by non-perturbative QCD effects, their mass $m_{a}$ and decay constant $f_{a}$ are independent parameters, potentially allowing stronger self-interactions.
If similar interactions were introduced between our scalar and SM fields—for instance, through a dimension-five operator of the form $g_{\phi\gamma\gamma}\phi F_{\mu\nu}\tilde{F}^{\mu\nu}$ or via terms $g_{\phi ee}\partial_{\mu}\phi\,\bar{e}\gamma^{\mu}\gamma_{5}e$—then similar astrophysical and cosmological limits derived for axions and ALPs with sub-eV masses could become relevant to this framework, even though the presence of strong repulsive self-interactions for SDM may change their effective applicability. The most significant constraints on axion DM in this mass range include (see Ref. [5] for a more comprehensive review):
[…]
  • Helioscope searches: For $m_{a}\lesssim 0.1~\mathrm{eV}$, the non-observation of solar axions (reconverted into X-ray photons via strong laboratory magnetic fields) implies limits onto the photon coupling of $g_{a\gamma\gamma}\lesssim 0.6\times 10^{-10}\,\mathrm{GeV}^{-1}$ (see, e.g., the CAST and IAXO experiments [22, 252, 23]).
  • Cosmological hot DM bounds: Thermalized axions with $m_{a}\gtrsim 0.7~\mathrm{eV}$ would contribute excessively to the relativistic energy density, violating CMB and large-scale structure constraints [107, 108, 339, 220, 458].
  • Laboratory experiments: Additional laboratory searches (including abracadabra-10 cm [367, 401, 366], SHAFT [219], and ADMX haloscopes [27, 163, 91, 45]) exclude complementary regions in the $(m_{a},g_{a\gamma\gamma})$ parameter space.
  • Gravitational waves: Couplings between ultralight DM and SM particles can induce oscillations in the fine-structure constant and electron mass, causing tiny periodic changes in interferometer components that modulate arm lengths and the measured gravitational-wave strain, allowing LIGO-Virgo-KAGRA to constrain such couplings [1].
In contrast, in the minimal version of the SDM model, these couplings are set to zero, leaving only gravitational interactions, thus implying that the model remains unconstrained by conventional searches for light, weakly coupled particles (but see discussion in Sec. 7 when considering condensate-induced long-range forces among baryons).
@@ 5.5 Vortices
$\vec{\nabla}\times\vec{v}=\frac{1}{m}2\pi\delta^{(3)}\big(\vec{r}-\vec{r}_{0}\big)\hat{z}\,$, (5.19)
@@
[…]
@@
[…]
@@
This naturally brings us to the point of vortex formation in our SDM halos, which may have some nonzero angular momentum, [^42] The presence of angular momentum in scalar DM models can also arise in an inhomogeneous density field, or if the patch has a nonspherical boundary, such as in Riemann S-ellipsoids [389]. see Refs. [413, 461, 97, 412, 260] for related works. Indeed, halos tend to form with rotation angular frequency parametrized as $\Omega\simeq\chi_{\rm s}\sqrt{G\rho}$, where the spin parameter $\chi_{\rm s}$ takes values in the range $0.01\lesssim\chi_{\rm s}\lesssim 0.1$, as inferred from N-body simulations [44].